OpenStudy (anonymous):

Express the quadratic equation 6/x-2(x+1)=1/4x-1/2 in the general form.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

multiply through out by the LCD

OpenStudy (anonymous):

LCD?

ganeshie8 (ganeshie8):

least common denominator

ganeshie8 (ganeshie8):

\[\large \dfrac{6}{x}-\dfrac{2}{x+1} = \dfrac{1}{4x} - \dfrac{1}{2}\] like this ?

OpenStudy (anonymous):